The mean amount spent by a family of four on food per month is $507 with a standard deviation of $73. Assuming that the food costs are normally distributed.

a. what is the probability that a family spends less than $406 per month? Give your answer to 4 decimals.

b. what is the probability that a family spends at least $711 per month? Give your answer to 4 decimals.
Answer #1

Solution :

Given that ,

mean = = $507

standard deviation = = $73

P(X<$406 ) = P[(X- ) / < (406 -507) /73 ]

= P(z < -1.38)

Using z table

= 0.0838

probability=0.0838

(b)

P(x >$711 ) = 1 - P(x<$711 )

= 1 - P[(x -) / < (711 -507) /73 ]

= 1 - P(z <2.79 )

Using z table

= 1 - 0.9974

= 0.0026

probability=0.0026